# Coupling Brownian loop soups and random walk loop soups at all polynomial scales

### v1

@@ Abstract
Lawler and Trujillo Ferreras [4] constructed a well-known coupling between the Brownian loop soups in $\mathbb{R}^{2}$ and the random walk loop soups on $\mathbb{Z}^{2}$ (one rescales the random walk loops by $1/N$, their time parametrizations by $1/(2N^{2})$, and let $N\to\infty$), which led to numerous applications. It nevertheless only holds for loops with time length at least $N^{\theta-2}$ for $\theta\in(2/3,2)$. In particular, there is no control on mesoscopic loops with time length less than $N^{-4/3}$ (i.e. roughly diameter less than $N^{-2/3}$).
In this paper, we find a simple way to remove the restriction $\theta>2/3$, so that such a coupling works for all $\theta\in(0,2)$, i.e. for loops at all polynomial scales. We also establish this coupling in any dimension $d\geq 1$ (i.e. for random walk loop soups on $\mathbb{Z}^{d}$ and Brownian loop soups on $\mathbb{R}^{d}$).
@@ 1.1 Lawler and Trujillo Ferreras’s coupling
$\mathcal{A}_{\lambda,N}=\{\Phi_{N}\gamma:\gamma\in\mathcal{A}_{\lambda}\}.$
@@
One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that for each $\lambda>0$, $\mathcal{A}_{\lambda}$ is a realization of the Brownian loop soup in $\mathbb{R}^{2}$, and $\widetilde{\mathcal{A}}_{\lambda}$ is a realization of the random walk loop soup on $\mathbb{Z}^{2}$. Both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$.
Moreover, for $\mathcal{A}_{\lambda,N}$ and $\widetilde{\mathcal{A}}_{\lambda,N}$ defined as above, there exists $c>0$ such that for every $r\geq 1$, $\lambda>0$, integer $N\geq 1$ and every $\theta\in(2/3,2)$, except perhaps on an event of probability at most $c(\lambda+1)r^{2}N^{2-3\theta}$, there is a one-to-one correspondence between the following two sets of loops
@@
$\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{-1}\log N.$
@@ 1.2 Main result
It is unclear at first sight whether $\theta>2/3$ can be removed, [^1] For example, this was believed to be imposssible in [1, Section 11.3]: “On the other hand, it is fairly clear from the proof of [4] that their result is sharp, and that the coupling described above cannot hold without the restriction $\theta>2/3$; that is, at all scales smaller than $N^{-2/3}$ some discrete and continuous loops somewhere will be quite different from one another.” since the value $2/3$ comes from a computation involving the following Taylor expansion of the total mass $q_{2}(n)$ of all random walk loops on $\mathbb{Z}^{2}$ with length $2n$ rooted at 0
@@
We will directly establish our result for random walks on $\mathbb{Z}^{d}$ for all $d\geq 1$. Note that the combinatorial formula (1.5) relies on a trick which is specific for $d=2$, namely we can rotate $\mathbb{Z}^{2}$ by $45$ degrees, so that the original simple random walk projected to the vertical and horizontal axes are independent simple random walks. For $d\geq 3$, it is more convenient to consider the continuous-time simple random walk, which spends an (independent) exponential time with mean $1$ at each vertex. We will rely on the following Taylor development of the total mas $q_{d}(t)$ of all continuous-time random walk loops on $\mathbb{Z}^{d}$ with time length $t$ rooted at 0 (see (2.3) and (2.4))
@@
Our proof requires only the first term in the Taylor series, unlike (1.5).
@@
$\chi_{N}(t)=\frac{k}{N^{2}}\quad\text{if}\quad\frac{a_{k}}{N^{2}}\leq t<\frac{a_{k+1}}{N^{2}}.$
@@
Fix $d\geq 1$. One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that for each $\lambda>0$,
  1. (1)
    $\mathcal{A}_{\lambda}$ is a realization of the Brownian loop soup in $\mathbb{R}^{d}$;
  2. (2)
    $\widetilde{\mathcal{A}}_{\lambda}$ is a realization of the continuous-time random walk loop soup on $\mathbb{Z}^{d}$ (or $\widetilde{\mathcal{A}}_{\lambda}$ can also be a discrete-time random walk loop soup if $d=1,2$).
Both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$. Moreover, for every $a>0$ and every $\theta\in(0,2)$, there exists $c>0$ such that for every $r\geq 1,\lambda>0$ and integer $N\geq 1$, except perhaps on an event of probability at most $c\lambda r^{d}N^{-a}$, there is a one-to-one correspondence between
@@
This strengthened coupling, which controls loops at any polynomial scale, rooted anywhere in a macroscopic domain, can have many potential applications. For example, in the two-dimensional case, it was proved by [6, 9] that the outer boundaries of the outermost clusters of the random walk loop soups with intensity $\lambda\in(0,1/2]$ converge to those of the Brownian loop soups (which are in turn distributed as conformal loop ensembles [8]). This convergence is nevertheless qualitative. If one wants to obtain a precise convergence rate, one may need to control the random walk loops at mesoscopic scales, everywhere in the domain simultaneously, since the outer boundary of a cluster is a global object.
On the other hand, for dimensions $d\geq 3$, to our knowledge, the convergence of random walk loop soups in $\mathbb{Z}^{d}$ to Brownian loop soups has not previously been written down, so we fill in this small gap in the literature. Note that it is clear that the same coupling also works for Brownian loop soups on the metric graph induced by $\mathbb{Z}^{d}$ (since its restriction to the vertices of $\mathbb{Z}^{d}$ is exactly a continuous-time random walk loop soup). This type of convergence has been used in e.g. [10, 7].
@@ 2.2 Discrete-time random walk loop soup
$\widetilde{q}_{d}(n):=|\nu_{n,z}|.$
For $d=1,2$, we have the exact combinatorial formula (1.5) for $\widetilde{q}_{2}(n)$, and
$\widetilde{q}_{1}(n)=(2n)^{-1}2^{-2n}\binom{2n}{n}$. (2.2)
@@ 2.3 Continuous-time random walk loop soup
Thanks to the following Taylor development of $p_{t}(0,0)$ (see e.g. [3, Theorem 2.13])
@@
2.4 KMT coupling
To couple a Brownian loop closely to a random walk loop, we rely on a variant of the classical coupling by Komlós, Major, and Tusnády [2]. Recall that the KMT coupling works for a wide class of random walks, and a comprehensive review of it can be found in [3, Chapter 7].
In particular, for any aperiodic irreducible walk $S_{n}=X_{1}+\cdots+X_{n}$ on $\mathbb{Z}$ such that there exist $\sigma,b>0$ with
$\mathbb{E}[X_{1}]=0,\quad\mathbb{E}[X_{1}^{2}]=\sigma^{2},\quad\mathbb{E}[\exp(b|X_{1}|)]<\infty,$ (2.6)
we can couple it with a Brownian motion (see e.g. [3, Theorem 7.1.1]).
@@ 2.4 KMT coupling
Suppose that $(S_{n})_{n\geq 0}$ is an aperiodic irreducible random walk on $\mathbb{Z}$ that satisfies (2.6). Then one can define on the same probability space a Brownian motion $B_{t}$ with variance $\sigma^{2}$ and $(S_{n})_{n\geq 0}$ such that the following holds. For each $\alpha<\infty$, there is a $c_{\alpha}$ such that
$\mathbb{P}\left[\max_{1\leq j\leq n}|S_{j}-B_{j}|\geq c_{\alpha}\log n\right]\leq c_{\alpha}n^{-\alpha}$
The idea of the proof of Theorem 2.1 has been used in [4] to prove the following coupling between a discrete-time simple random walk bridge on $\mathbb{Z}$ a Brownian bridge with length $2n$ (see [4, Theorem 6.4]). Note that even though the random walk bridges in Theorem 2.2 and Lemma 2.3 below are defined on discrete time and space, we still consider them as continuous curves parametrized by continuous time intervals, by linear interpolation. This is, however, different from continuous-time random walks.
@@
There exist $0<c,\alpha<\infty$, such that for every positive integer $n$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a simple random walk bridge $S$ on $\mathbb{Z}$, both with length $2n$, such that for all $r>0$,
By the trick of rotating $\mathbb{Z}^{2}$ by $45$ degrees (see the second paragraph of Section 1.2), the simple random walk on $\mathbb{Z}^{2}$ is the combination of two independent simple random walks on $\mathbb{Z}$, hence Theorem 2.2 implies the following coupling (see [4, Lemma 3.1] and the remark just below it).
@@
There exists a constant $c_{0}>0$, such that for each $\eta>0$, there exists $c(\eta)\in(0,\infty)$ so that the following holds. For every positive integer $n$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a two-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a two-dimensional simple random walk bridge $(S_{t},0\leq t\leq 2n)$ such that
$\mathbb{P}\left[\sup_{0\leq s\leq 1}\left|n^{-1/2}S_{2ns}-B_{s}\right|\geq c(\eta)n^{-1/2}\log n\right]\leq c_{0}n^{-\eta}.$
We will be interested in continuous-time random walk loops $\gamma=(S_{t},0\leq t\leq t_{\gamma})$ on $\mathbb{Z}^{d}$. Let $e_{1},\cdots,e_{d}$ be the $d$ unit vectors in $\mathbb{R}^{d}$. Then $S_{t}=S_{t/d}^{1}\,e_{1}+\cdots S^{d}_{t/d}\,e_{d}$ for all $0\leq t\leq t_{\gamma}$, where $S^{1}_{t},\cdots,S^{d}_{t}$ are independent continuous-time random walk bridges on $\mathbb{Z}$. Following the steps of [4], we can similarly deduce the following coupling (we will omit the proof since it is really analogous to that of Theorem 2.2).
Theorem 2.4.
There exist $0<c,\alpha<\infty$, such that for all $t>0$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a continuous-time simple random walk bridge $S$ on $\mathbb{Z}$, both with length $t$, such that for all $r>0$,
This then directly implies the following lemma.
Lemma 2.5.
Fix $d\geq 1$. There exists a constant $c_{0}(d)>0$, such that for each $\eta>0$, there exists $c(\eta,d)\in(0,\infty)$ so that the following holds. For every $t>0$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a $d$-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a $d$-dimensional continuous-time simple random walk bridge $(S_{s},0\leq s\leq t)$ such that
@@
3 Proof of the main theorem

### v2

@@ Abstract
Lawler and Trujillo Ferreras [6] constructed a well-known coupling between the Brownian loop soups in $\mathbb{R}^{2}$ and the random walk loop soups on $\mathbb{Z}^{2}$ (one rescales the random walk loops by $1/N$, their time parametrizations by $1/(2N^{2})$, and let $N\to\infty$), which led to numerous applications. It nevertheless only holds for loops with time length at least $N^{\theta-2}$ for $\theta\in(2/3,2)$. In particular, there is no control on mesoscopic loops with time length less than $N^{-4/3}$ (i.e. roughly diameter less than $N^{-2/3}$). This coupling was subsequently extended in [10] to $\mathbb{Z}^{d}$ with $d\geq 3$, for loops with time length at least $N^{\theta-2}$, for $\theta\in(2d/(d+4),2)$.
In this paper, we find a simple way to remove the restriction $\theta>2d/(d+4)$, so that such a coupling works for all $\theta\in(0,2)$, i.e. for loops at all polynomial scales. We establish couplings for both discrete-time and continuous-time random walk loop soups on $\mathbb{Z}^{d}$, for $d\geq 1$.
@@ 1.1 Lawler and Trujillo Ferreras’s coupling
$\mathcal{A}_{\lambda,N}=\{\Phi_{N}\gamma:\gamma\in\mathcal{A}_{\lambda}\}.$ (1.2)
@@
One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$. Moreover, for $\mathcal{A}_{\lambda,N}$ and $\widetilde{\mathcal{A}}_{\lambda,N}$ defined as above, there exists $c>0$ such that for every $r\geq 1$, $\lambda>0$, integer $N\geq 1$ and every $\theta\in(2/3,2)$, except perhaps on an event of probability at most $c(\lambda+1)r^{2}N^{2-3\theta}$, there is a one-to-one correspondence between the following two sets of loops
@@ 1.2 Main result
It is unclear at first sight whether $\theta>2/3$ can be removed, [^1] For example, this was believed to be impossible in [1, Section 11.3]: “On the other hand, it is fairly clear from the proof of [6] that their result is sharp, and that the coupling described above cannot hold without the restriction $\theta>2/3$; that is, at all scales smaller than $N^{-2/3}$ some discrete and continuous loops somewhere will be quite different from one another.” since the value $2/3$ comes from a computation involving the following Taylor expansion of the total mass $q_{2}(n)$ of all random walk loops on $\mathbb{Z}^{2}$ with length $2n$ rooted at 0
@@
Lawler and Trujillo Ferreras established this coupling in dimension $2$, but their proof has been extended to dimensions $d\geq 3$ by Sapozhnikov and Shiraishi in [10, Theorem 2.2], relying on the expansion (see e.g. [2])
$\widetilde{q}_{d}(n)=2\bigg(\frac{d}{4\pi n}\bigg)^{d/2}\bigg(1-\frac{d}{8n}+O\big(\frac{1}{n^{2}}\big)\bigg),$ (1.8)
where $\widetilde{q}_{d}(n)$ is the total mass of all random walk loops on $\mathbb{Z}^{d}$ with length $2n$ rooted at 0. Similarly, the coupling in [10] only works for loops with length greater than $N^{\theta-2}$, where
$2d/(d+4)<\theta<2,$ (1.9)
on an event with probability at least
$1-c(\lambda+1)r^{d}N^{-\min\left(\frac{d}{2},\;\theta\left(\frac{d}{2}+2\right)-d\right)}$ (1.10)
for some $c\in(0,\infty)$. They also obtained a slightly weaker bound $cN^{-1/4}\log N$ on the supreme distance between the coupled loops, instead of $cN^{-1}\log N$ in (1.6). This is because they established a weaker coupling between random walk bridges on $\mathbb{Z}^{d}$ with the Brownian bridges (see Lemma 2.3), for $d\geq 3$. Note that the strong KMT coupling in dimension $2$ established in [6] (see Lemma 2.2), as well as the exact combinatorial formula (1.7), relies on a trick which is specific to $d=2$: We can rotate $\mathbb{Z}^{2}$ by $45$ degrees, so that the original simple random walk projected to the vertical and horizontal axes are independent simple random walks.
In this paper, we consider both the discrete-time random walk loop soups (as in [6, 10]), and the continuous-time random walk loop soups on $\mathbb{Z}^{d}$, where every walk spends an (independent) exponential time with mean $1$ at each vertex. The latter version is more natural when one considers the occupation time field of the loop soup, due to a well-known isomorphism theorem [8]. In addition, the latter version is equal to the “restriction” of the Brownian loop soup on the cable graph of $\mathbb{Z}^{d}$ (i.e. we restrict the Brownian loops on the cable graph to stay on a vertex, before jumping to the next vertex). The two versions of random walk loop soups are the same, if one forgets the time parametrization and remove the single-vertex loops in the continuous-time random walk loop soup (see [5, Section 9]). However, since our coupling takes into account the time-parametrization of the loops, the couplings for the two versions of loop soups need separate proofs, which can be done in parallel (also see Remark 1.5). By abuse of notation, we use $\widetilde{\mathcal{A}}_{\lambda}$ to denote both versions of random walk loop soups. The actual version we use will be specified in the context.
The continuous-time random walk has the advantage that its movements in the $d$ coordinates are independent, hence we can obtain the same strong KMT coupling as in [6], yielding the same $cN^{-1}\log N$ bound as in (1.6) on the supreme distance between the coupled loops. In this case, we will rely on the following Taylor development of the total mas $q_{d}(t)$ of all continuous-time random walk loops on $\mathbb{Z}^{d}$ with time length $t$ rooted at 0 (see (2.3) and (2.4))
@@
Our proof requires only the first term in the Taylor series, unlike in [6, 10].
@@
$\chi_{N}(t)=\frac{k}{N^{2}}\quad\text{if}\quad\frac{a_{k}}{N^{2}}\leq t<\frac{a_{k+1}}{N^{2}}.$ (1.13)
@@
Fix $d\geq 1$. Let $\mathcal{A}_{\lambda}$ be a Brownian loop soup on $\mathbb{R}^{d}$, and $\widetilde{\mathcal{A}}_{\lambda}$ be a (continuous-time) random walk loop soup on $\mathbb{Z}^{d}$. One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$.
Moreover, for every $a>0$ and every $\theta\in(0,2)$, there exists $c>0$ such that for every $r\geq 1,\lambda>0$ and integer $N\geq 1$, except perhaps on an event of probability at most $c\lambda r^{d}N^{-a}$, there is a one-to-one correspondence between
@@
Note that Theorem 1.2 works for loops at any polynomial scale, rooted anywhere in a macroscopic domain. Moreover, the error probability (of the event on which the coupling fails) is at most $c\lambda r^{d}N^{-a}$, where $a$ can be made arbitrarily large, compared with Theorem 1.1 and (1.10). This strengthened coupling can have many potential applications. For example, in the two-dimensional case, it was proved by [9, 12] that the outer boundaries of the outermost clusters of the random walk loop soups with intensity $\lambda\in(0,1/2]$ converge to those of the Brownian loop soups (which are in turn distributed as conformal loop ensembles [11]). This convergence is nevertheless qualitative. If one wants to obtain a precise convergence rate, it then seems necessary to control the random walk loops at mesoscopic scales, everywhere in the domain simultaneously, since the outer boundary of a cluster is a global object.
The result for discrete-time random walk loop soups on $\mathbb{Z}^{d}$ is rather similar, but we still state it below in Theorem 1.3, so that the reader can directly compare it with [6] in the $d=2$ case (and with [10] in the $d\geq 3$ case). In the discrete-time case, we use a different sequence $\{a_{n}\}_{n\geq 1}$ defined in (3.14), which also satisfies (1.12). We keep the same definition (1.13) of $\chi_{N}$, but inputting the new sequence $\{a_{n}\}_{n\geq 1}$.
Theorem 1.3.
Fix $d\geq 1$. Let $\mathcal{A}_{\lambda}$ be a Brownian loop soup on $\mathbb{R}^{d}$, and $\widetilde{\mathcal{A}}_{\lambda}$ be a (discrete-time) random walk loop soup on $\mathbb{Z}^{d}$. One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$.
Moreover, for every $a>0$ and every $\theta\in(0,2)$, there exists $c>0$ such that for every $r\geq 1,\lambda>0$ and integer $N\geq 1$, except perhaps on an event of probability at most $c\lambda r^{d}N^{-a}$, there is a one-to-one correspondence between
$|t_{\gamma}-t_{\widetilde{\gamma}}|\leq c(d)N^{-2},\qquad$ (1.16)
$\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{-1}\log N\qquad\text{ if }d=1,2,$ (1.17)
$\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{\frac{a-d}{2d}}\log N\quad\;\text{ if }d\geq 3.$ (1.18)
Remark 1.4.
For $d\geq 3$, we get the bound $cN^{\frac{a-d}{2d}}\log N$ in (1.18), using the coupling between the random walk bridge and the Brownian bridge (Lemma 2.3) established in [10]. This bound tends to 0 as $N\to\infty$ only when $a<d$, hence we cannot reduce the error probability (of the event on which the coupling fails) below $c\lambda r^{d}N^{-d}$ in this case. Note that this does not prevent us from choosing $\theta$ close to 0, hence this coupling does still control loops at all polynomial scales. If we choose $a=d/2$, then we get the same $cN^{-1/4}\log N$ bound as in [10].
We believe that it is possible to improve the bound (1.18), by improving the coupling of Lemma 2.3. However, if we do not care about the time parametrization, for example if we only care about the Hausdorff distance between the coupled loops, then (1.15) already provides the strong bound $cN^{-1}\log N$.
Remark 1.5.
An immediate corollary of Theorems 1.2 and 1.3 is that the discrete-time and continuous-time random walk loop soups can also be coupled together, since they can both be coupled with the Brownian loop soup. This coupling is, however, quite different from the obvious coupling where we let the two versions of loop soups have the same trace on $\mathbb{Z}^{d}$ (except for the single-vertex loops). Note that two random walk loops with the same trace, respectively parametrized by discrete and continuous time, typically do not have time lengths close to each other in the sense of (1.16).
@@ 2 Preliminaries
In this section, we recall the definitions and some basic properties of the Brownian loop soup, the discrete-time and continuous-time random walk loop soups on $\mathbb{Z}^{d}$, and record several versions of couplings between Brownian bridges and random walk bridges.
@@ 2.2 Discrete-time random walk loop soup
$\widetilde{q}_{d}(n):=|\nu_{n,z}|.$ (2.2)
The combinatorial formula of $\widetilde{q}_{2}(n)$ is given in (1.7) (and a similar formula can easily be obtained for $\widetilde{q}_{1}(n)$). In general, the Taylor development of $\widetilde{q}_{d}(n)$ is given by (1.8).
@@ 2.3 Continuous-time random walk loop soup
Thanks to the following Taylor development of $p_{t}(0,0)$ (see e.g. [5, Theorem 2.13])
@@
2.4 Coupling Brownian bridges and random walk bridges
In order to couple a Brownian bridge closely to a random walk bridge on the same probability space, we rely on several variants of the classical coupling by Komlós, Major, and Tusnády [4]. We will consider discrete-time random walk bridges in Theorem 2.1, Lemmas 2.2 and 2.3, and continuous-time random walk bridges in Theorem 2.4 and Lemma 2.5. For both cases, we regard the random walk bridges as continuous curves, by linear interpolation.
First of all, we need the following coupling between a discrete-time simple random walk bridge on $\mathbb{Z}$ and a Brownian bridge with length $2n$, established in [6, Theorem 6.4].
@@ 2.4 Coupling Brownian bridges and random walk bridges
There exist $0<c,\alpha<\infty$, such that for every positive integer $n$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a (discrete-time) random walk bridge $S$ on $\mathbb{Z}$, both with length $2n$, such that for all $r>0$,
By the trick of rotating $\mathbb{Z}^{2}$ by $45$ degrees (see the second paragraph of Section 1.2), the simple random walk on $\mathbb{Z}^{2}$ is the combination of two independent simple random walks on $\mathbb{Z}$, hence Theorem 2.1 implies the following coupling (see [6, Lemma 3.1] and the remark just below it).
Lemma 2.2.
There exists a constant $c_{0}>0$, such that for each $\eta>0$, there exists $c(\eta)\in(0,\infty)$ so that the following holds. For every positive integer $n$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a two-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a two-dimensional (discrete-time) random walk bridge $(S_{t},0\leq t\leq 2n)$ such that
$\mathbb{P}\left[\sup_{0\leq s\leq 1}\left|n^{-1/2}S_{2ns}-B_{s}\right|\geq c(\eta)n^{-1/2}\log n\right]\leq c_{0}n^{-\eta}.$ (2.6)
In [10, Lemma 4.2], the authors established the following coupling between random walk bridges on $\mathbb{Z}^{d}$ and $d$-dimensional Brownian bridges, for $d\geq 3$. They got a weaker $n^{-1/4}$ factor in (2.7), compared with $n^{-1/2}$ in (2.6), due to the technical difficulty arising from the fact that the movements of the random walk in $d$ coordinates are not independent.
Lemma 2.3.
Fix $d\geq 3$. For any $\eta>0$, there exists $c(\eta,d)\in(0,\infty)$ so that the following holds. For every positive integer $n$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a $d$-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a (discrete-time) random walk bridge $(S_{t},0\leq t\leq 2n)$ on $\mathbb{Z}^{d}$ such that
$\mathbb{P}\left[\sup_{0\leq s\leq 1}\big|(2n/d)^{-1/2}S_{2ns}-B_{s}\big|>c(\eta,d)n^{-1/4}\log n\right]\leq c(\eta,d)n^{-\eta}.$ (2.7)
We now turn to the continuous-time random walk bridge on $\mathbb{Z}$. Its coupling with the Brownian bridge is analogous to the discrete-time case, and has been proved in [3], where the authors actually considered random bridges constructed through a rather general class of random walks with i.i.d. jumps. Let us state the following version that we will use.
@@
There exist $0<c,\alpha<\infty$, such that for all $t>0$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a (continuous-time) random walk bridge $S$ on $\mathbb{Z}$, both with length $t$, such that for all $r>0$,
We now consider a continuous-time random walk bridge $\gamma=(S_{t},0\leq t\leq t_{\gamma})$ on $\mathbb{Z}^{d}$. Let $e_{1},\cdots,e_{d}$ be the $d$ unit vectors in $\mathbb{R}^{d}$. Then $S_{t}=S_{t/d}^{1}\,e_{1}+\cdots S^{d}_{t/d}\,e_{d}$ for all $0\leq t\leq t_{\gamma}$, where $S^{1}_{t},\cdots,S^{d}_{t}$ are independent continuous-time random walk bridges on $\mathbb{Z}$. Theorem 2.4 then directly implies the following lemma.
@@
Fix $d\geq 1$. There exists a constant $c_{0}(d)>0$, such that for each $\eta>0$, there exists $c(\eta,d)\in(0,\infty)$ so that the following holds. For every $t>0$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a $d$-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a $d$-dimensional (continuous-time) random walk bridge $(S_{s},0\leq s\leq t)$ such that
@@
3 Proof of the main theorems